OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies each set of conditions. Passes through (-9,-3), perpendicular to y=-5/3-8

OpenStudy (the_fizicx99):

The perp of -5/3 is 3/5 Y-y1=M(x-x1) Y-(-3))=3/5(x-(-9))

OpenStudy (the_fizicx99):

Solve dat

OpenStudy (anonymous):

y=3/5x+6 ?

OpenStudy (gamer56):

ok so first tell me what you thimk you should...when you respond to me tag my name

OpenStudy (anonymous):

so now what do I do ? is this right so far y=3/5x+6

OpenStudy (gamer56):

looks correct to me.

OpenStudy (anonymous):

now do I have to solve for b by putting in this formula y=mx+b ?

OpenStudy (anonymous):

Someone help please!!!!

OpenStudy (ranga):

y - (-3) = 3/5 * (x - (-9)) y + 3 = 3/5x + 27/5 y = 3/5x + 27/5 - 3 y = 3/5x + (27 - 15) / 5 y = 3/5x + 12/5 is the equation of the line in slope intercept form.

OpenStudy (the_fizicx99):

Sorry I was afk >.<

OpenStudy (anonymous):

Thanks ranga!